### v1

@@ IV Mean-Field Game
With the AoI calculations in the above subsection, we have provided a complete formulation of the $N$-user game problem. A suitable solution concept for the above game is that of Nash equilibrium [24]. However, its computation becomes intractable due to the high population regime. Thus, to alleviate this issue, we design Nash policies with the additional attractive feature that each agent uses only its local information. In this regard, we leverage the framework of MFGs [7]. Under the latter, we consider the limiting case ($N=\infty$) of the finite-agent system (called the MF system). In this scenario, the individual agent’s deviations from equilibrium policies become insignificant due to the presence of infinite number of agents. Hence, it suffices to consider the viewpoint of generic agents (representing the population of a specific type) which play against a mass distribution rather than each individual agent. This then allows for the computation of the MF equilibrium, which constitutes an optimal policy of a generic agent which is consistent with that of the population.
In the next subsection, we set up the generic agent’s optimization problem using the MFG framework.
IV-A Generic Device’s Optimization Problem
In this subsection, we state the optimization problem faced by the generic device in the infinite population regime. The packets arrive at a generic device of type $\phi$ at mean rate $\lambda_{\phi}$. The tasks are split by employing an i.i.d. Bernoulli distributed random variable with a mean $p_{\phi}$ into two independent Poisson processes with respective means $\lambda_{\phi}p_{\phi}$ and $\lambda_{\phi}\bar{p}_{\phi}$. Further, the mean service rates of the generic transmitter and the generic local processor are given as $\mu_{1,\phi}$ and $\mu_{2,\phi}$, respectively.
Let us define the average AoI of the packets of class 1 as ${\Delta}_{\phi}(p_{\phi},\mu_{1\phi},\mu_{2\phi},\rho):=\lim_{N\rightarrow\infty}\Delta^{(N)}_{i}(\bm{p},\bm{\mu}_{1},\bm{\mu}_{2})$ where $\rho:=\lim_{N\rightarrow\infty}\rho^{(N)}$. The latter exists, for instance, when the service rate of the ES increases proportionally to the number of devices, i.e., when $\mu_{3}^{(N)}=N\mu_{3}$, where $\mu_{3}>0$. The term $\rho$ serves as the MF approximation to the coupling term $\rho^{(N)}$ in the finite-device system and can be viewed as the mean load on the ES in the infinite-device system. Then, we have the generic device optimization problem as follows.
@@ IV-A Generic Device’s Optimization Problem
where
$J_{\rho}(p_{\phi},\mu_{1\phi},\mu_{2\phi})$ $\!:=\!\mu_{1,\phi}+\eta\mu_{2\phi}^{3}\!+\!V{\Delta}_{\phi}(p_{\phi},\mu_{1\phi},\mu_{2\phi},\rho)$ (8)
and ${\Delta}_{\phi}(p_{\phi},\mu_{1\phi},\mu_{2\phi},\rho)$ is given in (6a).
@@
Briefly, for a given value of $\rho$, the generic agent solves for an optimal policy using the optimality condition. Consequently, it uses the obtained policy to regenerate $\rho$ using the consistency condition. The mean-field equilibrium (MFE), which constitutes the pair $((p_{\phi,\text{MFE}},\mu_{1\phi,\text{MFE}},\mu_{2\phi,\text{MFE}}),\rho_{\text{MFE}})$, is then given by the fixed point of the composite map induced by 1) and 2) for all $\phi\in\Phi$. Detailed fixed point iteration process is given in Algorithm 1 below.
@@
9:Output: $\hat{\rho}^{(m)},(\hat{p}^{(m)}_{\phi},\hat{\mu}_{1\phi}^{(m)},\hat{\mu}_{2\phi}^{(m)}),~{}\forall\phi.$
@@ V Numerical Results
Here, we provide a numerical computation of the MFE for a population of a single type. In the first numerical study, in Fig. 4, we observe that as the mean loading at the ES increases (on the $x$-axis), the optimal probability of using the local processor (on the $y$-axis) increases and that of offloading to the ES decreases. This should be expected since if the ES is heavily loaded, the device is better-off serving tasks locally to incur a lower AoI. Next, in Fig. 5, we plot the variation of the MFE as a function of the arrival rate $\lambda$ and service rate $\mu_{3}$. From the figure, we observe that at equilibrium increasing arrival rate offloads more computations, thereby increasing the ES loading. On the other hand, increasing ES service rate increases offloading by the device, but with a decrease in the mean ES loading, thereby suggesting a slower than linear optimal rate of task offloading by the devices.
@@ VI Discussion and Conclusion
As a recap, in this work we have considered a timely task computation problem in a MEC system where the devices can either process their tasks on their local processors or offload them to an ES. We have developed a finite-agent Nash game and a MFG model for the task offloading problem in MEC systems, and numerically computed an equilibrium solution to the MF system. In the future, we plan to investigate two important aspects of the above developments: (1) Characterization of the conditions ensuring existence and (possible) uniqueness of the MFE, and (2) a theoretical analysis of the performance of the MFE solution on the original finite-agent system.

### v2

@@ IV Mean-Field Game
With the AoI calculations in the above subsection, we have provided a complete formulation of the $N$-user game problem. A suitable solution concept for the above game is that of seeking a Nash equilibrium policy [29], i.e., a policy from which no user can deviate to receive a lower cost. However, its computation becomes intractable due to the high population regime. Thus, to alleviate this issue, we design Nash policies with the additional attractive feature that each user uses only its local policy information. In this regard, we leverage the framework of MFGs [8]. Under the latter, we consider the limiting case ($N=\infty$) of the finite-user system (called the MF system). In this scenario, the individual user’s deviations from equilibrium policies become insignificant due to the presence of infinite number of users. Hence, it suffices to consider the viewpoint of generic user (representing the population of a specific type) which play against a mass distribution rather than each individual user. This then allows for the computation of the MF equilibrium in a completely decentralized manner, which constitutes an optimal policy of a generic user which is consistent with that of the population. With the above prelude, let us set up the MFG as follows.
Consider a generic device of type $\phi$ in the infinite population regime. The packets arrive at the device at mean rate $\lambda_{\phi}$. The tasks are split by employing an i.i.d. Bernoulli distributed random variable with a mean $p_{\phi}$ into two independent Poisson processes with respective means $\lambda_{\phi}p_{\phi}$ and $\lambda_{\phi}\bar{p}_{\phi}$. Further, the mean service rates of the generic transmitter and the generic local processor are given as $\mu_{1,\phi}$ and $\mu_{2,\phi}$, respectively. Let us define the average AoI of the packets of class 1 as ${\Delta}_{\phi}(p_{\phi},\mu_{1\phi},\mu_{2\phi},\rho):=\lim_{N\rightarrow\infty}\Delta^{(N)}_{i}(\bm{p},\bm{\mu}_{1},\bm{\mu}_{2})$ where we define $\rho^{(N)}:=\frac{\lambda_{e}^{(N)}}{\mu^{(N)}_{3}}$, and $\rho:=\lim_{N\rightarrow\infty}\rho^{(N)}$. The latter exists, for instance, when the service rate of the ES increases proportionally to the number of devices, i.e., when $\mu_{3}^{(N)}=N\mu_{3}$ for $\mu_{3}>0$, which is also what we consider for numerical evaluation purposes. The term $\rho$ serves as the MF approximation to the coupling term $\rho^{(N)}$ in the finite-device system and can be viewed as the mean load on the ES in the infinite-device system. Then, we have the generic device optimization problem as follows.
@@
where $J_{\rho}(p_{\phi},\mu_{1\phi},\mu_{2\phi})\!:=\!t_{L_{\phi}}\mu_{1\phi}+t_{T_{\phi}}\eta\mu_{2\phi}^{3}\!+\!V{\Delta}_{\phi}(p_{\phi},\mu_{1\phi},\mu_{2\phi},\rho)$, ${\Delta}_{\phi}(p_{\phi},\mu_{1\phi},\mu_{2\phi},\rho)$ is given in (6a) and $t_{L_{\phi}},t_{T_{\phi}}$ denote the busy periods of $L_{\phi}$ and $T_{\phi}$, respectively, of the generic device.
@@
Briefly, for a given value of $\rho$, the generic user solves for an optimal policy using the optimality condition. Consequently, it uses the obtained policy to regenerate $\rho$ using the consistency condition. The mean-field equilibrium (MFE), which constitutes the pair $((p_{\phi,\text{MFE}},\mu_{1\phi,\text{MFE}},\mu_{2\phi,\text{MFE}}),\rho_{\text{MFE}})$, is then given by the fixed point of the composite map induced by 1) and 2) for all $\phi\in\Phi$. Detailed fixed point iteration process is given in Algorithm 1 below.
@@
9:Output: Last iterate: $\hat{\rho}^{(m)},(\hat{p}^{(m)}_{\phi},\hat{\mu}_{1\phi}^{(m)},\hat{\mu}_{2\phi}^{(m)}),~{}\forall\phi.$
@@ V Numerical Results
Here, we provide a numerical computation of the MFE for a population of a single type. In the first numerical study, in Fig. 5, (for $V=10,\eta=5,\lambda=2.5,P_{T,max}=1,f_{max}=0.3$) we observe that as the mean loading at the ES increases (on the $x$-axis), the optimal probability of using the local processor (on the $y$-axis) increases and that of offloading to the ES decreases. This should be expected since if the ES is heavily loaded, the device is better-off serving tasks locally to incur a lower AoI. Next, in Fig. 6, (for $V=10,\eta=0.5,P_{T,max}=1,f_{max}=0.3$) we plot the variation of the MFE as a function of the arrival rate $\lambda$ and service rate $\mu_{3}$. From the figure, we observe that at equilibrium increasing arrival rate offloads more computations, thereby increasing the ES loading. On the other hand, increasing ES service rate increases offloading by the device, but with a decrease in the mean ES loading, thereby suggesting a slower than linear optimal rate of task offloading by the devices.
@@ VI Discussion and Conclusion
As a recap, in this work we have considered a timely task computation problem in a dense-user MEC system where the devices can either process their tasks on their local processors or offload them to an ES. We have developed a finite-user Nash game and a MFG model for the task offloading problem in MEC systems, and provided a low complexity algorithm to compute decentralized equilibrium solutions to the MF system. In the future, we plan to investigate the theoretical aspects of the above developments, particularly related to the characterization of the conditions ensuring existence and (possible) uniqueness of the MFE in ultra-dense user networks.